Find the surface area of the function x= ^3 square root of y between x=0 and x=1 when revolved around the axis


1
Expert's answer
2021-05-24T04:11:56-0400

Solution:

"x=\sqrt[3]{y}
\\\Rightarrow y=x^3=f(x)"

Between "x=0,x=1" , revolved around "x"-axis.

Surface area"=2\pi\int_a^bf(x)\sqrt{1+[f'(x)]^2}dx"

"=2\pi\int_0^1 x^3\sqrt{1+[3x^2]^2}dx
\\=2\pi\int_0^1 x^3\sqrt{1+9x^4}dx\ ...(i)"

Put "1+9x^4=t"

"\Rightarrow 36x^3=\frac{dt}{dx}
\\\Rightarrow x^3 dx=\frac{dt}{36}"

When "x=0,t=1"

When "x=1,t=10"

So, (i) becomes,

Surface area"=\frac{2\pi}{36}\int_1^{10}\sqrt{t}dt"

"=\frac{\pi}{18}[\frac23t^{3/2}]_1^{10}
\\=\frac{\pi}{27}[{10}^{3/2}-1]
\\=3.563\ units^2"